# Remote Sensing Spatio-Temporal Vision-Language Models: A Comprehensive Survey

### v2

@@ IV-C LLM-driven Task-Specific Spatio-Temporal VLMs
TABLE IV: Some representative studies based on LLMs for remote sensing spatio-temporal vision-language understanding.
Spatio-Temporal VLMs Method Release Time Visual Encoder LLM Fine-tuning Code Link
LLM-driven Task-Specific Spatio-Temporal VLMs PromptCC [72] 2023.10 CLIP-ViT-B/32 GPT-2 Prompt Tuning [Link]
@@
Unified Spatio-Temporal Vision-Language Foundation Model Change-Agent [101] 2024.03 Segformer Chatgpt Frozen [Link]
@@
EarthDial [115] 2024.12 InternViT-300M Phi-3-mini Fully Fine-tuning
@@
IV-D Unified Spatio-Temporal Vision-Language Foundation Models
Recent advancements in artificial intelligence have led to the rise of foundation models—large-scale, pre-trained models designed to be general-purpose and adaptable across a wide range of downstream tasks. Trained on massive and diverse datasets, these models demonstrate strong generalization capabilities and can be efficiently adapted to new tasks with minimal task-specific fine-tuning. In the field of remote sensing, foundation models have garnered increasing attention across different domains [187]. These include image perception foundation models (e.g., RingMo [188], SpectralGPT [189], HyperSIGMA [190], and MTP [191]), image generation foundation models (e.g., DiffusionSat [192], Text2Earth [161], and MetaEarth [193]), and vision-language foundation models (e.g., RSGPT [194], GeoChat [195], VHM [196], LHRS-Bot [197], EarthGPT [198], and EarthMarker [199]).
In the context of remote sensing spatio-temporal vision-language understanding, recent research has begun to explore unified spatio-temporal vision-language foundation models, which aim to integrate diverse spatio-temporal tasks into a single, coherent framework [200]. Unlike conventional approaches that develop separate models for tasks such as change captioning, change question answering, or change detection, unified spatio-temporal models leverage LLMs to interpret user instructions and generate task-specific outputs in an autoregressive manner. Typically, they combine vision encoders for processing remote sensing temporal imagery with language decoders for generating responses, creating a versatile and reusable pipeline. These unified models offer significant advantages in terms of generalization, scalability, and interactive capability. By aligning spatio-temporal visual features with the language modality and supporting multi-task instruction tuning, they pave the way for flexible and extensible remote sensing systems that can seamlessly adapt to diverse application scenarios.
@@ IV-D1 Unified Data Representation
A fundamental challenge in building unified foundation models is how to represent heterogeneous task outputs—ranging from bounding boxes to segmentation masks—in a language-friendly format that enables a large language model (LLM) to generate them autoregressively as plain text. This unified textual formulation not only removes the need for task-specific heads but also facilitates joint multi-task training and instruction tuning across various remote sensing applications.
Bounding boxes are typically categorized into two formats: horizontal bounding boxes (HBB) and oriented bounding boxes (OBB). Both are normalized into a fixed range (e.g., [0, 100]) and represented as sequences of numeric tokens. The HBB format describes a box as $[x_{min},y_{min},x_{max},y_{max}]$, where $(x_{min},y_{min})$ and $(x_{max},y_{max})$ denote the top-left and bottom-right corners, respectively. This representation is widely adopted in models such as TEOChat [68], EarthGPT [198], and SkyEyeGPT [201]. To express orientation, the OBB format introduces an additional rotation angle $\theta$. There are several ways to express OBBs: 1) Corner-point form: GeoRSMLLM [128] and Falcon [127] opt for a more geometry-preserving representation by encoding the four corner points directly as $[(x_{1},y_{1}),(x_{2},y_{2}),(x_{3},y_{3}),(x_{4},y_{4})]$. 2) Center-based form: SkysenseGPT [202] represents boxes as $[c_{x},c_{y},w,h,\theta]$, where $(c_{x},c_{y})$ is the center coordinate, w and h are the box width and height, and $\theta$ is the rotation angle. 3) Hybrid form: EarthDial [115] and GeoChat [195] adopt a variation combining both HBB coordinates and an orientation term, such as $[x_{min},y_{min},x_{max},y_{max},\theta]$. These representations enable the LLM to output localization information for detection or grounding tasks purely in text. Notably, HBB is a special case of OBB when $\theta$ = 0, which allows unified training under a unified format.
Pixel-wise Masks represent dense outputs and are common in change detection tasks. They must also be converted into language-compatible formats. Several methods encode masks as: 1) Semantic descriptor sequences, where masks are transformed into structured textual descriptions that reflect spatial layout, as in RSUniVLM [114] and Text4Seg [203]; 2) Polygon-based encodings, where object contours are represented as ordered point sets [(x1, y1), …, (xn, yn)], as in GeoRSMLLM [128] and Falcon [127].
@@
Figure 11: A representative approach for Unified Spatio-Temporal Vision-Language Foundation Model driven by LLMs. TEOChat  [68] supports analyzing image sequences of arbitrary temporal lengths and can perform multiple tasks, including temporal scene classification, change detection, and temporal image question answering.
@@ IV-D2 Representative Methods
A pioneering example is TeoChat[68]. As illustrated in Fig. 11, TeoChat accepts multi-temporal images of arbitrary length and encodes them as visual tokens, which are used as prefix inputs to a pre-trained Llama2 [126]. By unifying various task outputs in natural language form and employing joint training over multiple objectives, TeoChat can follow user instructions and simultaneously supports multiple tasks, such as temporal scene classification, change detection, change referring expressions, and temporal image question answering. This design reflects a clear transition from isolated, single-task pipelines toward end-to-end, instruction-driven multimodal foundation models with strong generalization across spatio-temporal tasks.
@@
These unified spatio-temporal VLMs mark a crucial step toward general-purpose remote sensing vision-language models, capable of executing multiple tasks flexibly under natural language instructions. They not only simplify deployment in real-world scenarios but also push the boundary of what multimodal models can achieve in the remote sensing spatio-temporal domain.
@@
Change-Agent [101] 2024.03 Spatio-Temporal Change Interpretation [Link]
@@ IV-E LLM-driven Remote Sensing Vision-Language Agents
In the field of remote sensing, LLM-based agents are beginning to show unique potential [204, 101, 205, 207, 215, 206]. Table IV summarizes representative methods. For instance, RS-Agent [205] consists of a central LLM-based controller, a retrieval-augmented module, and a toolset of remote sensing models. The controller parses user instructions and plans task workflows, while the retrieval module supplements the LLM with relevant domain knowledge or solutions from external databases to facilitate choosing appropriate tools. The toolset includes diverse vision models for tasks such as classification and change detection. Depending on task complexity, RS-Agent can either invoke a single tool or compose multiple models in a sequential pipeline to accomplish multi-step reasoning objectives. Li et al. provide a survey on remote sensing agents in [216].
In the context of remote sensing temporal image understanding, Change-Agent [101] is a representative example tailored for interpreting remote sensing temporal visual changes. It adopts an LLM as the central reasoning engine and integrates specialized modules (e.g., change detection and captioning) as visual tools, further augmented by external utilities. Upon receiving a complex instruction—such as “count the number of changed houses”—Change-Agent first decomposes the task, then invokes a change detection model to generate change masks, and finally generates and executes code (e.g., Python scripts) to post-process the masks. This process not only produces accurate and verifiable outputs, but also alleviates hallucination problems (e.g., object counting) often seen in end-to-end unified vision-language models. Furthermore, the Change-Agent supports customized outputs, such as detecting specific change categories (e.g., vegetation loss or road expansion), offering flexible solutions for diverse real-world needs.
@@ V Evaluation Metrics and Dataset
To rigorously assess the performance of spatio-temporal vision–language models (TVLMs), both quantitative metrics and representative datasets are indispensable. In this section, we first review the principal evaluation metrics—grouped by output modality—and then present the major datasets, organized according to their annotation schemas.
@@ V-A3 Localization Metrics
When models output spatial regions (e.g., bounding boxes, masks) revealing regions of interest within images, localization metrics quantify how well predictions spatially align with ground truth:
@@
    $\text{IoU}_{i}=\frac{\mid\text{A}_{i}\cap\text{B}_{i}\mid}{\mid\text{A}_{i}\cup\text{B}_{i}\mid}$
    $\text{MIoU}=\frac{1}{\text{N}}\sum_{i=1}^{\text{N}}\text{IoU}_{i}$
    Where N is the number of samples, $\text{A}_{i}$ and $\text{B}_{i}$ are the predicted region and ground truth of the $i$-th sample.
  • CIoU: Cumulative Intersection over Union (CIoU) [222] metric evaluates the intersection over union ratio between the predicted regions and the ground truth across multiple IoU thresholds. It provides a more comprehensive evaluation, especially in complex scenarios with large prediction deviations or overlapping objects.
    $\text{CIoU}=\frac{1}{\text{T}}\sum_{i=1}^{\text{T}}\text{IoU}_{i}$
    Where T is the number of selected IoU thresholds, $\text{A}_{i}$ and $\text{B}_{i}$ represent the predicted region and the ground truth region at the $i$-th IoU threshold, respectively.
Figure 13: Spatio-temporal vision-language datasets are typically built upon existing change detection datasets or unlabeled temporal remote sensing images. The annotations are created using manual labeling, rule-based methods, or with the assistance of ChatGPT. Based on their annotation formats, we categorize existing datasets into three categories: 1) Dataset Matching Temporal Images and Text. 2) Dataset Matching Temporal Images, Text and Mask. 3) Dataset Matching Temporal Images, Instruction and Response.
V-B Remote Sensing Spatio-Temporal Vision-Language Datasets
The development of remote sensing spatio-temporal vision–language models relies critically on high-quality datasets. These datasets provide critical supervision for diverse tasks such as change captioning, change question answering, and change grounding. As shown in Fig. 13, spatio-temporal vision-language datasets are typically built upon existing change detection datasets or unlabeled temporal remote sensing images. The annotations are created using manual labeling, rule-based methods, or with the assistance of ChatGPT. Based on their annotation formats, we categorize existing datasets into three three categories: datasets pairing temporal images with text, datasets incorporating both textual and pixel-level mask annotations, and datasets offering instruction and response.
@@ V-B1 Dataset Matching Temporal Images and Text
Dataset Time Image Size Image Resolution Image Pairs Captions Masks Temporal Image Data Source Anno. Link
DUBAI CCD [70] 2022.08 50$\times$50 30m 500 2,500 - Landsat-7 imagery Manual [Link]
LEVIR CCD [70] 2022.08 256$\times$256 0.5m 500 2,500 - LEVIR-CD [48] Manual [Link]
LEVIR-CC [26] 2022.11 256$\times$256 0.5m 10,077 50,385 - LEVIR-CD [48] Manual [Link]
@@
LEVIR-MCI [101] 2024.03 256$\times$256 0.5m 10,077 50,385 building, road LEVIR-CC [26] Manual [Link]
LEVIR-CDC [106] 2024.11 256$\times$256 0.5m 10,077 50,385 building LEVIR-CC [26] Manual [Link]
WHU-CDC [106] 2024.11 256$\times$256 0.075m 7,434 37,170 building WHU-CD [230] Manual [Link]
SECOND-CC [93] 2025.01 256$\times$256 0.3$\sim$3m 6,041 30,205 6 classes SECOND [229] Manual [Link]
@@
  • LEVIR-CC [26]: Compared to previous small datasets (i.e., DUBAI CCD and LEVIR CCD), LEVIR-CC is a large-scale change captioning benchmark dataset widely used in current research. It includes 10,077 bi-image pairs (256×256 pixels, 0.5m/pixel). The images are sourced from the LEVIR-CD dataset [48] and cover 20 regions across Texas, USA, spanning 5–14 years. Each pair is annotated with five human-written descriptions, totalling 50,385 diverse captions, with an average sentence length of 11 words. The dataset was curated to exclude trivial changes (e.g., lighting variations) and emphasize substantial changes like the appearance or disappearance of ground objects (e.g., buildings, roads, and vegetation).
@@ V-B2 Dataset Matching Temporal Images, Text, and Masks
These datasets facilitate spatio-temporal vision-language learning by simultaneously providing textual descriptions and pixel-level change masks. They are crucial for developing multi-task learning across both spatial localization and semantic understanding, and exploring the complementary relationship between change detection and change captioning. Comparative dataset statistics are provided in Table VI, with some examples shown in Fig. 13.
@@
TEOChatlas [68] 2024.10 554,071 - 1$\sim$8 xBD [233], S2Looking [234]), QFabric [235], fMoW [232] Auto. [Link]
EarthDial [115] 2024.12 11.11 Million - 1$\sim$4 fMoW [232], TreeSatAI-Time-Series [236], MUDS [237], xBD [233], QuakeSet [238] Manual & Auto. [Link]
@@
Falcon_SFT [127] 2025.03 78 Million 5.6 Million 1$\sim$2 CDD [240], EGY-BCD [241], HRSCD [242], LEVIR-CD [48], MSBC [243], MSOSCD [243], NJDS [244], S2Looking [234], SYSU-CD[228], WHU-CD [230] Auto. [Link]

### v3

@@ IV-C LLM-driven Task-Specific SpatioTemporal VLMs
TABLE IV: Some representative studies based on LLMs for remote sensing spatiotemporal vision-language understanding.
RS-STVLMs Method Time Visual Encoder LLM Fine-tuning Code Link
LLM-driven Task-Specific SpatioTemporal VLMs PromptCC [98] 2023.10 CLIP-ViT-B/32 GPT-2 Prompt Tuning [Link]
@@
Unified SpatioTemporal Vision-Language Foundation Model Change-Agent [127] 2024.03 Segformer Chatgpt Frozen [Link]
@@
EarthDial [141] 2024.12 InternViT-300M Phi-3-mini Fully Fine-tuning [Link]
@@
DVLChat  [143] 2025.05 SAM Qwen2.5-VL LoRA
Figure 10: A representative approach for LLM-driven task-specific spatiotemporal VLMs. PromptCC  [98] first introduces LLMs into the change captioning task.
IV-D Unified SpatioTemporal Vision-Language Foundation Models
Recent advancements in artificial intelligence have led to the rise of foundation models, large-scale pre-trained models designed to be general-purpose and adaptable across a wide range of downstream tasks. Trained on massive and diverse datasets, these models demonstrate strong generalization capabilities and can be efficiently adapted to new tasks with minimal task-specific fine-tuning. In the field of remote sensing, foundation models have garnered increasing attention across different domains [39]. These include image perception foundation models (e.g., RingMo [244], SpectralGPT [245], HyperSIGMA [246], MTP [247] and DynamicVis [248]), image generation foundation models (e.g., DiffusionSat [249], Text2Earth [204], and MetaEarth [250]), and vision-language foundation models (e.g., RSGPT [33], GeoChat [34], VHM [251], LHRS-Bot [36], EarthGPT [37], EarthMarker [252], and GeoLLaVA-8K [253]).
In the context of remote sensing spatiotemporal vision-language understanding, recent research has begun to explore unified spatiotemporal vision-language foundation models, which aim to integrate diverse spatiotemporal tasks into a single, coherent framework [254]. Unlike conventional approaches that develop separate models for tasks such as change captioning, change question answering, or change detection, unified spatiotemporal models leverage LLMs to interpret user instructions and generate task-specific outputs in an autoregressive manner. Typically, they combine vision encoders for processing remote sensing temporal imagery with language decoders for generating responses, creating a versatile and reusable pipeline. These unified models offer significant advantages in terms of generalization, scalability, and interactive capability. By aligning spatiotemporal visual features with the language modality and supporting multi-task instruction tuning, they pave the way for flexible and extensible remote sensing systems that can seamlessly adapt to diverse application scenarios.
@@ IV-D1 Unified Data Representation
A fundamental challenge in building unified foundation models is how to represent heterogeneous task outputs (i.e., bounding boxes, segmentation masks) in a language-friendly format that enables an LLM to generate them autoregressively as plain text. This unified textual formulation not only removes the need for task-specific heads but also facilitates joint multi-task training and instruction tuning across various remote sensing applications.
Bounding boxes are typically categorized into two formats: horizontal bounding boxes (HBB) and oriented bounding boxes (OBB). Both are normalized into a fixed range (e.g., [0, 100]) and represented as sequences of numeric tokens. The HBB format describes a box as $x_{min},y_{min},x_{max},y_{max}]$, where $(x_{min},y_{min})$ and $(x_{max},y_{max})$ denote the top-left and bottom-right corners, respectively. This representation is widely adopted in models such as TEOChat [94], EarthGPT [37], and SkyEyeGPT [255]. To express orientation, the OBB format introduces an additional rotation angle $\theta$. There are several ways to express OBBs: 1) Corner-point form: GeoRSMLLM [155] and Falcon [154] opt for a more geometry-preserving representation by encoding the four corner points directly as $(x_{1},y_{1}),(x_{2},y_{2}),(x_{3},y_{3}),(x_{4},y_{4})]$. 2) Center-based form: SkysenseGPT [256] represents boxes as $c_{x},c_{y},w,h,\theta]$, where $(c_{x},c_{y})$ is the center coordinate, w and h are the box width and height, and $\theta$ is the rotation angle. 3) Hybrid form: EarthDial [141] and GeoChat [34] adopt a variation combining both HBB coordinates and an orientation term, such as $x_{min},y_{min},x_{max},y_{max},\theta]$. These representations enable the LLM to output localization information for detection or grounding tasks purely in text. Notably, HBB is a special case of OBB when $\theta$ = 0, which allows unified training under a unified format.
Pixel-wise Masks represent dense outputs and are common in change detection tasks. They must also be converted into language-compatible formats. Several methods encode masks as: 1) Semantic descriptor sequences, where masks are transformed into structured textual descriptions that reflect spatial layout, as in RSUniVLM [140] and Text4Seg [257]; 2) Polygon-based encodings, where object contours are represented as ordered point sets $(x_{1},y_{1}),...,(x_{n},y_{n})]$, as in GeoRSMLLM [155] and Falcon [154].
@@
Figure 11: A representative approach for Unified SpatioTemporal Vision-Language Foundation Model driven by LLMs. TEOChat  [94] supports analyzing image sequences of arbitrary temporal lengths and can perform multiple tasks, including temporal scene classification, change detection, and temporal image question answering.
@@ IV-D2 Representative Methods
A pioneering example is TeoChat[94]. As illustrated in Fig. 11, TeoChat accepts multi-temporal images of arbitrary length and encodes them as visual tokens, which are used as prefix inputs to a pre-trained Llama2 [153]. By unifying various task outputs in natural language form and employing joint training over multiple objectives, TeoChat can follow user instructions and simultaneously supports multiple tasks, such as temporal scene classification, change detection, change referring expressions, and temporal image question answering. This design reflects a clear transition from isolated, single-task pipelines toward end-to-end, instruction-driven multimodal foundation models with strong generalization across spatiotemporal tasks.
@@
These unified spatiotemporal VLMs mark a crucial step toward general-purpose remote sensing vision-language models, capable of executing multiple tasks flexibly under natural language instructions. They not only simplify deployment in real-world scenarios but also push the boundary of what multimodal models can achieve in the remote sensing spatiotemporal domain.
@@
Change-Agent [127] 2024.03 SpatioTemporal Change Interpretation [Link]
@@ IV-E LLM-driven Remote Sensing Vision-Language Agents
In the field of remote sensing, LLM-based agents are beginning to show unique potential [258, 127, 259, 261, 269, 260]. Table V summarizes representative methods. For instance, RS-Agent [259] consists of a central LLM-based controller, a retrieval-augmented module, and a toolset of remote sensing models. The controller parses user instructions and plans task workflows, while the retrieval module supplements the LLM with relevant domain knowledge or solutions from external databases to facilitate choosing appropriate tools. The toolset includes diverse vision models for tasks such as classification and change detection. Depending on task complexity, RS-Agent can either invoke a single tool or compose multiple models in a sequential pipeline to accomplish multi-step reasoning objectives. Li et al. provide a survey on remote sensing agents in [270].
In the context of remote sensing temporal image understanding, Change-Agent [127] is a representative example tailored for interpreting remote sensing temporal visual changes. It adopts an LLM as the central reasoning engine and integrates specialized modules (e.g., change detection and captioning) as visual tools, further augmented by external utilities. Upon receiving a complex instruction, such as “count the number of changed houses”, Change-Agent first decomposes the task, then invokes a change detection model to generate change masks, and finally generates and executes code (e.g., Python scripts) to post-process the masks. This process not only produces accurate and verifiable outputs, but also alleviates hallucination problems (e.g., object counting) often seen in end-to-end unified vision-language models. Furthermore, the Change-Agent supports customized outputs, such as detecting specific change categories (e.g., vegetation loss or road expansion), offering flexible solutions for diverse real-world needs.
@@ V Evaluation Metrics and Dataset
To rigorously assess the performance of RS-STVLMs, both quantitative metrics and representative datasets are indispensable. In this section, we first review the principal evaluation metrics grouped by output modality, and then present datasets organized according to their annotation schemas.
@@ V-A3 Localization Metrics
When models output spatial regions (e.g., masks, bounding boxes) revealing regions of interest within images, localization metrics quantify how well predictions spatially align with ground truth:
@@
    $\text{IoU}_{k}=\frac{\mid\text{P}_{k}\cap\text{G}_{k}\mid}{\mid\text{P}_{k}\cup\text{G}_{k}\mid}$
    $\text{MIoU}=\frac{1}{\text{K}}\sum_{k=1}^{\text{K}}\text{IoU}_{k}$
    Where K is the total number of semantic classes, $\text{P}_{i}$ and $\text{G}_{i}$ are the predicted region and ground truth of the $i$-th semantic class.
  • cIoU: The cumulative Intersection over Union (cIoU) [277] metric evaluates the overall intersection over union ratio between the predicted regions and the ground truth at the dataset level. cIoU aggregates the total intersection and union across all samples before computing their ratio. This can prevent individual samples with small regions from disproportionately influencing the overall score. The cIoU is computed as:
    $\text{cIoU}=\frac{\sum_{i=1}^{N}|P_{i}\cap G_{i}|}{\sum_{i=1}^{N}|P_{i}\cup G_{i}|}$
    Where N is the total number of samples in the dataset, $\text{P}_{i}$ and $\text{G}_{i}$ are the predicted region and ground truth of the $i$-th sample.
  • F1 Score: The F1 score is the harmonic mean of Precision and Recall, and it reflects the balance between false positives and false negatives. F1 is more sensitive to prediction correctness and is commonly used in tasks with class imbalance. It can be computed per class $k$ as:
    $\text{F1}_{k}=\frac{2\cdot TP_{k}}{2\cdot TP_{k}+FP_{k}+FN_{k}}$
    The mean F1 score across all classes is given by:
[…]
    Where K is the total number of classes, and $TP_{k},FP_{k},FN_{k}$ represent the true positives, false positives, and false negatives for class $k$, respectively.
Figure 13: Spatiotemporal vision-language datasets are typically built upon existing change detection datasets or unlabeled temporal remote sensing images. The annotations are created using manual labeling, rule-based methods, or with the assistance of ChatGPT. Based on their annotation formats, we categorize existing datasets into three categories: 1) Dataset Matching Temporal Images and Text. 2) Dataset Matching Temporal Images, Text and Mask. 3) Dataset Matching Temporal Images, Instruction and Response.
V-B Remote Sensing SpatioTemporal Vision-Language Datasets
The development of RS-STVLMs relies critically on high-quality datasets. These datasets provide critical supervision for diverse tasks such as change captioning, change question answering, and change grounding. As shown in Fig. 13, spatiotemporal vision-language datasets are typically built upon existing change detection datasets or unlabeled temporal remote sensing images. The annotations are created using manual labeling, rule-based methods, or with the assistance of ChatGPT. Based on their annotation formats, we categorize existing datasets into three three categories: datasets pairing temporal images with text, datasets incorporating both textual and pixel-level mask annotations, and datasets offering instruction and response.
@@ V-B1 Dataset Matching Temporal Images and Text
Dataset Time Image Size Image Resolution Image Pairs Texts Masks Temporal Image Data Source Anno. Link
DUBAI CCD [96] 2022.08 $\times50$ 30m 500 2,500 - Landsat-7 imagery Manual [Link]
LEVIR CCD [96] 2022.08 $\times256$ 0.5m 500 2,500 - LEVIR-CD [71] Manual [Link]
LEVIR-CC [28] 2022.11 $\times256$ 0.5m 10,077 50,385 - LEVIR-CD [71] Manual [Link]
@@
[…]
LEVIR-MCI [127] 2024.03 $\times256$ 0.5m 10,077 50,385 building & road LEVIR-CC [28] Manual [Link]
LEVIR-CDC [132] 2024.11 $\times256$ 0.5m 10,077 50,385 building LEVIR-CC [28] Manual [Link]
WHU-CDC [132] 2024.11 $\times256$ 0.075m 7,434 37,170 building WHU-CD [286] Manual [Link]
[…]
@@
  • LEVIR-CC [28]: Compared to previous small datasets (i.e., DUBAI CCD and LEVIR CCD), LEVIR-CC is a large-scale change captioning benchmark dataset widely used in current research. It includes 10,077 bi-image pairs (256×256 pixels, 0.5m/pixel). The images are sourced from the LEVIR-CD dataset [71] and cover 20 regions across Texas, USA, spanning 5-14 years. Each pair is annotated with five human-written descriptions, totalling 50,385 diverse captions, with an average sentence length of 11 words. The dataset was curated to exclude trivial changes (e.g., lighting variations) and emphasize substantial changes like the appearance or disappearance of ground objects (e.g., buildings, roads, and vegetation).
@@ V-B2 Dataset Matching Temporal Images, Text, and Masks
These datasets facilitate spatiotemporal vision-language learning by simultaneously providing textual descriptions and pixel-level change masks. They are crucial for developing multi-task learning across both spatial localization and semantic understanding, and exploring the complementary relationship between change detection and change captioning. Comparative dataset statistics are provided in Table VI, with some examples shown in Fig. 13.
@@
[…]
DVL-Suite [143] 2025.05 69,926 15,063 6.9 (Average) U.S. National Agriculture Imagery Program (NAIP) Manual & Auto.